OpenStudy (anonymous):

How do I evaluate log927? Could someone be so kind as to explain how to do this?? I do not have these notes, and I am studying for a test.

OpenStudy (ash2326):

@cayssaday do you know the basic formula for logarithm?

OpenStudy (anonymous):

Unfortunately not :(

OpenStudy (ash2326):

Ok, I will explain : Suppose you have \[a=\log_b c\] It implies \[b^a=c\] Suppose you have \[2=\log_{10} 100\] it implies \[10^2=100\] Do you follow?

OpenStudy (anonymous):

Yes!

OpenStudy (ash2326):

ok there are some properties of logarithm, you should remember these always \[\log_a b+\log_a c=\log_a {b\times c}\] I forgot to mention a is called the base, so if you have \[\log_2 4=> \text{2 is the base here}\] \[\log_2 {10} => \text{10 is the base here}\] do you follow?

OpenStudy (anonymous):

Yes I do!

OpenStudy (anonymous):

Wait but why is the base of the first one 2 and the second one 10? Shouldn't the second one be 2 as well?

OpenStudy (ash2326):

Base can be anything, base should be greater than 0 but not 1 \[base>0 , base\ne 1\]

OpenStudy (anonymous):

Oh okay cool! Makes sense!

OpenStudy (ash2326):

ok, can you find the value of this? \[\log _2 8=?\]

OpenStudy (anonymous):

I guess I do not quite understand ...

OpenStudy (ash2326):

Say the value is x \[\log_2 8=x\] it implies \[8=2^x\] what will be the power of 2 which is equal to 8?

OpenStudy (anonymous):

would it be 4? Oh gosh I am sorry I really am not good at math.

OpenStudy (ash2326):

2x 2 x 2=?

OpenStudy (anonymous):

6

OpenStudy (ash2326):

try again? \[2\times 2 \times 2=?\]

OpenStudy (anonymous):

8 lol

OpenStudy (ash2326):

what power of 2 is 8?

OpenStudy (anonymous):

ummm 3?

OpenStudy (ash2326):

good, so \[\log_2 8=\text {3}\]

OpenStudy (ash2326):

do you follow?

OpenStudy (anonymous):

Yes I am sorry for the trouble thank you so much!!!

OpenStudy (ash2326):

No problem, let's work on your problem \[\log _9{27}\] what's the base here?

OpenStudy (anonymous):

9

OpenStudy (ash2326):

ok, we will need to find a number which when raised to the power of 9 is 27 Do you have any idea?

OpenStudy (anonymous):

Well.. 9 ^1 = 9 and 9 ^2 = 81.... But 9*3=27...

OpenStudy (whpalmer4):

uh, no, 9 is raised to a number...

OpenStudy (whpalmer4):

Here's a question: is 27 a power of 3?

OpenStudy (anonymous):

Is it 1.5?

OpenStudy (ash2326):

oops @whpalmer4 Cayssaday 9 x 9 x 9=729 not 27 Do you know the square of 27?

OpenStudy (whpalmer4):

yes, it is !

OpenStudy (ash2326):

@cayssaday how did you find that it's 1.5?

OpenStudy (whpalmer4):

I was wondering the same thing :-)

OpenStudy (anonymous):

the square root of 27 is 5.19 right?

OpenStudy (ash2326):

What is the square of 27? \[27^2=?\]

OpenStudy (whpalmer4):

yes, it is. \[\sqrt{27} = \sqrt{3*9} = 3*\sqrt{3} \approx 3*1.7320508 = 5.196...\]

OpenStudy (anonymous):

(9)^x =(9)^(1/2) = 3

OpenStudy (ash2326):

Godd Cayssaday, keep going

OpenStudy (ash2326):

*good

OpenStudy (anonymous):

Idk i just guessed sort of

OpenStudy (ash2326):

your approach is correct. According to the log property of addition \[\log_9(27)=\log_9(3\times 9)=\log_93+\log_99\] I think you can validate your guess now, can't you?

OpenStudy (whpalmer4):

\(5*5 + 2*5*0.2 + 0.2*0.2 = 25 + 2 +0.04 \approx 27\) if we approximate 5.196^2 by \((5+0.2)^2\) Well, if you know that \[\sqrt{9} = 3\]and\[3^3=27\]then you can go directly to \[(\sqrt{9})^3 = 27\]or\[((9)^{1/2})^3 = 27\]or\[9^{3/2} = 27\]

OpenStudy (whpalmer4):

@jim_thompson5910 probably has an interesting technique up his sleeve as well..

jimthompson5910 (jim_thompson5910):

Here's how you do it using the change of base formula and rule # 3 on this link http://www.purplemath.com/modules/logrules.htm \[\Large y = \log_{9}(27)\] \[\Large y = \frac{\log(27)}{\log(9)} \ ... \text{Using Change Of Base Formula}\] \[\Large y = \frac{\log(3^3)}{\log(3^2)}\] \[\Large y = \frac{3*\log(3)}{2*\log(3)} \ ... \text{Using Rule 3}\] \[\Large y = \frac{3*\cancel{\log(3)}}{2*\cancel{\log(3)}}\] \[\Large y = \frac{3}{2}\] \[\Large y = 1.5\] Side Note: the change of base formula is explained here http://www.purplemath.com/modules/logrules5.htm

OpenStudy (the_fizicx99):

There's really nothing more than knowing they're multiples of each other. I used basic math and got 1.5 .-.

OpenStudy (ash2326):

@cayssaday Do you follow ?

OpenStudy (anonymous):

Yes sorry I am taking so long to reply, I am taking notes. I do online school and there isn't really lessons or anything so It's really challenging. You guys are helping SOO much thank you!!!

OpenStudy (anonymous):

Evaluate log25125 The answer is 1.5 right?

OpenStudy (whpalmer4):

Yes, because \(\sqrt{25} = 25^{1/2} = 5\) and \(25^1 = 25\), so \(125 = 25*5\) and \(\log_{25} 125 = \log_{25} (25*5) = \log_{25} 25 + \log_{25} 5 = 1+1/2\) Just like before...

OpenStudy (anonymous):

Yes yes awesome it is making sense :D